# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v9

@@ 4.3 Theorem on Model Generalization Error Estimation
$p_{M}(x)\begin{cases}\approx p_{q}(x),&x\in S_{ou}\cap S_{oq}\\ =(1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|),&x\in S_{oq}\backslash S_{ou}\end{cases}$ (4.12)
Using KL divergence to measure the discrepancy between the true distribution and the hypothesized model distribution, we have Goodfellow2016 :
$D_{KL}(p_{q}(x)\|p_{M}(x))=$ $\sum_{x\in S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ (4.13)
$=$ $\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$
According to (4.12), in the overlapping region:
$\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}\to\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{q}(x)}=0$ (4.14)
@@
$=-H(p_{q})-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))$
$=-H(p_{q})-\log((1-P_{Mq})/(|S_{oq}|-|S_{ou}\cap S_{oq}|))\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)$ (1)
$=(1-P_{Mq})\log((|S_{oq}|-|S_{ou}\cap S_{oq}|)/(1-P_{Mq}))-H(p_{q})$ (4.15)
Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their total variation distance. By Pinsker’s inequality Pinsker1964 , we have:
$\text{TVD}(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\|p_{M}(x))}$
$=\sqrt{\frac{1}{2}(\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)})}$
$\approx\sqrt{\frac{1}{2}((1-P_{Mq})\log((|S_{oq}|-|S_{ou}\cap S_{oq}|)/(1-P_{Mq}))-H(p_{q}))}$ (4.16)
@@
Since we have already assumed that the carrier states $S_{M}(M,Tu_{m})$ and $S_{cq}(cq,Tq_{m})$ of $I_{u}=\langle ou,Tu_{h},S_{ou},M,Tu_{m},S_{M}\rangle$ and $I_{q}=\langle oq,Tq_{h},S_{oq},cq,Tq_{m},S_{cq}\rangle$ are both represent the physical realizations of $S_{ou}$ and $S_{oq}$ in model M and its interface, respectively. It follows that equation (4.11) characterizes the upper bound of the model’s TVD. This completes the proof.
Unlike traditional generalization theories such as VC dimension or Rademacher complexity, Theorem 6, based on the MLT-MF framework, provides an explicit quantitative estimate for the model’s TVD. This estimate is grounded in the overlap between the test and training data $P_{Mq}$ and $|S_{oq}|-|S_{ou}\cap S_{oq}|$, as well as the entropy $H(p_{q})$ of the test data distribution over the non-overlapping region. These metrics serve as indicators of how much the testing data shares information with the model’s prior knowledge and how unpredictable the unseen portion is.
Because TVD characterizes the alignment between distributions at a fundamental level—independent of any particular model architecture or task—it is generally applicable to the analysis of model training performance. In fact, for classification tasks with 0-1 loss, the generalization error upper bound is equivalent to the TVD Rodríguez2021 . For regression tasks, the upper bound on generalization error is given by the product of TVD and the Lipschitz constant of the loss function Devroye2013 . Thus, TVD is a meaningful and informative measure for estimating model generalization error.
This shows that Theorem 6 offers a new perspective and practical guidance for data collection, model design, and training for construction. In particular, Theorem 6 reveals that as long as the problem space is contained within the model’s information space, the model’s TVD can be made arbitrarily close to zero in a statistical sense. This is a fundamental result in machine learning:
Corollary 2 (TVD for Ideal Models) If the model’s information encompasses all relevant information for the problem, then the TVD for these problems is approximately zero.
@@ 5 Experimental Results
Using the feedforward neural network model listed in the appendix, we conducted regression experiments on two bivariate root functions: $y_{1}=\sqrt{4x_{1}^{2}+3x_{2}^{2}}$ and $y_{2}=\sqrt{2x_{1}^{2}+x_{2}^{2}}$. Both the training and test datasets were generated randomly. We evaluated the models’ prediction errors under various conditions, assessed the impact and computational cost of ethical and safety mechanisms, and explored the challenge of providing multiple types of explanations for the same output. These experiments serve to illustrate and validate the definitions and theorems presented in Section 4.
5.1 Prediction Error Experiments
To evaluate the model’s predictive performance in both overlapping and non-overlapping cases between the training and test datasets, we used the following data setup: 10,000 training samples were generated with $x_{1}$ and $x_{2}$ independently and uniformly distributed over $[0.5,1.5]$, and 10,000 test samples were drawnuniformly from within a circle of radius 1.0 centered at (1.5, 1.0). The model’s learning rate was set to $\eta=0.01$.
Figure 2 illustrates the experimental conditions and outcomes:
Figure 2: Model Prediction Error Experiments and Results Under Different Conditions. (a) The blue rectangular region represents the training dataset comprising 10,000 samples, while the red circular region represents the test dataset, also with 10,000 samples. Among these, 4,951 samples overlap with the training set. (b) For each of the training steps (10, 50, 100, 200, 500, 1000, 4000), the variation in root mean square error (RMSE) is shown for 10 randomly selected test samples from the training dataset. (c) In the non-overlapping region between the test and training datasets, the RMSEs for all 5,049 test samples are compared between the conservative uniform prediction strategy and the TVD upper bound. (d) For 10 randomly selected test samples from the non-overlapping region, the changes in RMSE are shown for training steps (10, 30, 300, 600, 1000, 4000, 7000), along with comparisons to both the TVD upper bound and the RMSE from the conservative uniform prediction.
@@ 5.1 Prediction Error Experiments
(c) For all test samples from the non-overlapping region, a conservative uniform prediction strategy was used. For almost all samples (except a few outliers), their RMSE is less than $\text{TVD}=\sqrt{((1-0.4951)\ln((10000-4951)/(1-0.4951))-2.9741)/2}\approx 0.9155$. This indicates that even in regression tasks, although TVD is not strictly equivalent to generalization error, it can still effectively capture the distributional characteristics of generalization error.

### v10

@@ 5.3 Theorem on TVD Upper Bounds
[…]
$p_{M}(x)=\begin{cases}\approx p_{q}(x),&\text{if }x\in S_{ou}\cap S_{oq}\\[3.00003pt] \frac{1-P_{Mq}}{|S_{oq}|-|S_{ou}\cap S_{oq}|},&\text{if }x\in S_{oq}\backslash S_{ou}\end{cases}$ (5.12)
Using KL-divergence to measure the difference between the true distribution and the hypothesized model distribution, we have (Kullback and Leibler, 1951):
$D_{KL}(p_{q}(x)\parallel p_{M}(x))$ $=\sum_{x\in S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$
$=\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ (5.13)
According to (5.12), in the overlapping region,
$\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}$ $\approx\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{q}(x)}$
$=0$ (5.14)
@@
$=-H(p_{q})-\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{(1-P_{Mq})}{|S_{oq}|-|S_{ou}\cap S_{oq}|}$
$=-H(p_{q})-\log\frac{(1-P_{Mq})}{|S_{oq}|-|S_{ou}\cap S_{oq}|}\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)$
$=(1-P_{Mq})\log\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}}-H(p_{q})$ (5.15)
Since both $p_{M}(x)$ and $p_{q}(x)$ are probability distributions defined on the same measurable space, we can use KL-divergence to estimate an upper bound on their total variation distance. By Pinsker’s inequality (Pinsker, 1964), we have: …
$TVD(p_{q},p_{M})$ $\leq\sqrt{\frac{1}{2}D_{KL}(p_{q}(x)\parallel p_{M}(x))}$
$=\sqrt{\frac{1}{2}\left(\sum_{x\in S_{ou}\cap S_{oq}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}+\sum_{x\in S_{oq}\backslash S_{ou}}p_{q}(x)\log\frac{p_{q}(x)}{p_{M}(x)}\right)}$
$\approx\sqrt{\frac{1}{2}\left((1-P_{Mq})\log\frac{|S_{oq}|-|S_{ou}\cap S_{oq}|}{1-P_{Mq}}-H(p_{q})\right)}$ (5.16)
@@
Since we have already assumed that the carrier states $S_{M}(M,{Tu}_{m})$ and $S_{cq}(cq,{Tq}_{m})$ of $Iu=\left\langle ou,{Tu}_{h},S_{ou},M,{Tu}_{m},S_{M}\right\rangle$ and $Iq=\left\langle oq,{Tq}_{h},S_{oq},cq,{Tq}_{m},S_{cq}\right\rangle$ are both represent the physical realizations of $S_{ou}$ and $S_{oq}$ in model $M$ and its interface, respectively. It follows that equation (5.11) characterizes the upper bound of the model’s TVD. This completes the proof.
Distinct from traditional generalization theories such as VC dimension and Rademacher complexity, Theorem 3, based on the MLT-MF framework, utilizes $P_{Mq}$ and $\left|S_{oq}\right|-\left|S_{ou}\cap S_{oq}\right|$—which reflect the overlap between test data and the model’s original information—as well as the entropy $H\left(p_{q}\right)$ of test data in the non-overlapping region, as evaluation metrics to provide an explicit quantitative estimation of the TVD. Since TVD reflects the degree of distributional alignment of the model, independent of specific models and tasks, it is generally applicable for analyzing training performance. Moreover, for classification tasks with 0-1 loss, the upper bound of the generalization error is equivalent to the TVD (Devroye et al., 2013), while for regression tasks, the upper bound of the generalization error is given by the product of the TVD and the Lipschitz loss (Devroye et al., 2013). Therefore, TVD serves as an indicator for estimating a model’s generalization error. In this way, Theorem 3 offers a novel perspective and practical guidance for model design and training data acquisition.
In general, input information is inevitably affected by noise, and arbitrary predictive strategies can be adopted in regions not covered by training data. According to Definition 3 on structural decomposition of noisy information, we can thus obtain more general results regarding the model’s TVD.
Theorem 4 (TVD Upper Bound for Arbitrary Prediction Strategies with Noisy Inputs and Uncovered Training Regions) Let the base space $X$ be a finite set. The training information domain of the model $M$ is $S_{ou}\subseteq X$. The true problem domain is $S_{oq}\subseteq X$, and the true distribution $p_{q}$ satisfies $\sum_{x\in S_{oq}}{p_{q}(x)}=1$. The noisy information ontological state $\SoqSoq+\widetilde{S_{oq}}=S_{oq}\backslash$ $\cup S_{oq}^{+}$, as defined in Definition 3, replaces $S_{oq}$ as the actual input to the model $M$. The noise loss probability mass is $P_{l}=\sum_{x\in S_{oq}^{-}}{p_{q}(x)}\in[0,1]$, and the noise addition probability mass is $P_{a}=\sum_{x\in S_{oq}^{+}}{p^{+}(x)}\in[0,1]$, where $p^{+}$ is a non-negative kernel on the noise addition region, satisfying $1-P_{l}+P_{a}>0$. The distribution of the noisy input $\widetilde{S_{oq}}$ is given by $\widetilde{p_{q}}(x)=(p_{q}(x)\bullet 1\left\{x\notin S_{oq}^{-}\right\}+p^{+}(x)\bullet 1\left\{x\in S_{oq}^{+}\right\})/(1-P_{l}+P_{a})$. Let $V=\widetilde{S_{oq}}\backslash S_{ou}$ denote the region not covered by training information, with uncovered probability mass $P_{uncov}=\sum_{x\in V}{\widetilde{p_{q}}(x)}\in[0,1]$. The conditional distribution on $V$ is $r(x)=\widetilde{p_{q}}(x)/P_{uncov}$ (and when $P_{uncov}=0$, all terms involving $V$ below are taken as 0). The predictive strategy $p_{0}$ on the uncovered region $V$ is an arbitrary distribution, satisfying $\sum_{x\in V}{p_{0}(x)=1,}p_{0}(x)\geq 0$. The model distribution $p_{M}$ is determined by training on the covered region $U=\widetilde{S_{oq}}\cap S_{ou}$, and on $V$, $p_{M}(x)=P_{uncov}p_{0}(x)$. Assume that the model's fitting error on the covered region is $E_{fit}=D_{KL}\left(\widetilde{p_{q}}|_{U}\parallel p_{M}|_{U}\right)$, where $\widetilde{p_{q}}|_{U}$ and $p_{M}|_{U}$ denote the distributions restricted to $U$ and normalized (under the ideal conditions assumed in Theorem 3, $E_{fit}$ tends to zero).
Thus, the upper bound of the TVD for model $M$ with respect to noisy input information is given by $TVD\left(\widetilde{p_{q}},p_{M}\right)\leq\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}}$ (5.17)
The upper bound of the TVD for model $M$ with respect to the true information is
$TVD\left(p_{q},p_{M}\right)\leq\left\{\begin{array}[]{r}P_{l}+\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}},\ \ \ \ \ \ \ \ \ \ \ if\ P_{l}\geq P_{a},\\ \frac{P_{a}}{1-P_{l}+P_{a}}+\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}},\ \ \ \ \ \ if\ P_{l}<P_{a}.\end{array}\right$.\ (5.18)
Proof: Based on the conditions provided in the theorem, we have:
$D_{KL}\left(\widetilde{p_{q}}\parallel p_{M}\right)=\sum_{x\in\widetilde{S_{oq}}}{\widetilde{p_{q}}(x)\log\frac{\widetilde{p_{q}}(x)}{p_{M}(x)}}$
$\sum_{x\in U}{\widetilde{p_{q}}(x)\log\frac{\widetilde{p_{q}}(x)}{p_{M}(x)}}+\sum_{x\in V}{\widetilde{p_{q}}(x)\log\frac{\widetilde{p_{q}}(x)}{p_{M}(x)}}$
$D_{KL}\left(\widetilde{p_{q}}|_{U}\parallel p_{M}|_{U}\right)+\sum_{x\in V}{P_{uncov}r(x)\log\frac{P_{uncov}r(x)}{P_{uncov}p_{0}(x)}}$
$E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)$ (5.17)
Here, $E_{fit}$ reflects the degree of fit of $M$ on the region covered by the training information, and its probabilistic upper bound can be characterized by measures such as VC dimension, Rademacher complexity, or training stability. Moreover, by applying Pinsker’s inequality, the first conclusion of the theorem can be obtained:
$TVD\left(\widetilde{p_{q}},p_{M}\right)\leq\sqrt{\frac{D_{KL}\left(\widetilde{p_{q}}\parallel p_{M}\right)}{2}}=\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}}$ (5.18)
Since the distribution of the noisy input $\widetilde{S_{oq}}$ is
$\widetilde{p_{q}}(x)=\frac{p_{q}(x)\bullet 1\left\{x\notin S_{oq}^{-}\right\}+p^{+}(x)\bullet 1\left\{x\in S_{oq}^{+}\right\}}{1-P_{l}+P_{a}}$ (5.19)
It follows that
$TVD\left(p_{q},\widetilde{p_{q}}\right)=(\sum_{x\in X}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|)/2$
$(\sum_{x\in S_{oq}^{-}}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|+\sum_{x\in S_{oq}^{+}}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|+\sum_{x\in X\backslash(S_{oq}^{-}\cup S_{oq}^{+})}\left|p_{q}(x)-\widetilde{p_{q}}(x)\right|)/2$
$(\sum_{x\in S_{oq}^{-}}\left|p_{q}(x)\right|+\sum_{x\in S_{oq}^{+}}\left|p^{+}(x)/(1-P_{l}+P_{a})\right|+\sum_{x\in X\backslash(S_{oq}^{-}\cup S_{oq}^{+})}\left|p_{q}(x)-p_{q}(x)/(1-P_{l}+P_{a})\right|)/2$
$(P_{l}+\frac{P_{a}}{1-P_{l}+P_{a}}+\frac{\left|P_{l}-P_{a}\right|\sum_{x\in X\backslash(S_{oq}^{-}\cup S_{oq}^{+})}{p_{q}(x)}}{1-P_{l}+P_{a}})/2=(P_{l}+\frac{P_{a}}{1-P_{l}+P_{a}}+\frac{\left|P_{l}-P_{a}\right|(1-P_{l})}{1-P_{l}+P_{a}})/2$
$\left\{\begin{array}[]{r}P_{l},\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ if\ P_{l}\geq P_{a},\\ \frac{P_{a}}{1-P_{l}+P_{a}},\ \ \ \ \ \ \ \ \ \ \ \ \ \ if\ P_{l}<P_{a}.\end{array}\right$.\ (5.22)
By applying the triangle inequality, we have
$TVD\left(p_{q},p_{M}\right)\leq TVD\left(p_{q},\widetilde{p_{q}}\right)+TVD\left(\widetilde{p_{q}},p_{M}\right)$
$\leq\left\{\begin{array}[]{r}P_{l}+\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}},\ \ \ \ \ \ \ \ if\ P_{l}\geq P_{a},\\ \frac{P_{a}}{1-P_{l}+P_{a}}+\sqrt{\frac{\left[E_{fit}+P_{uncov}D_{KL}\left(r\parallel p_{0}\right)\right]}{2}},\ if\ P_{l}<P_{a}.\end{array}\right$.\ (5.23)
Thus, the second conclusion of the theorem is established, completing the proof.
Theorem 4 unifies the noise structure, coverage geometry, and uncovered-region strategies within a computable TVD upper bound framework. Specifically, the TVD upper bounds for model $M$ with respect to both noisy actual inputs and true inputs are expressed in terms of the noise loss mass $P_{l}$, addition mass $P_{a}$, fitting error over the covered region $E_{fit}$, uncovered mass $P_{uncov}$, and the KL divergence $D_{KL}\left(r\parallel p_{0}\right)$ between the conditional distribution on the uncovered region and the prediction strategy. This formulation reveals several important insights: the loss and addition masses determine the deviation between the true target and the noisy target; the fit over the covered region, the coverage structure, and the divergence between the conditional distribution and prediction strategy on the uncovered region jointly determine the model's deviation from the noisy target. Consequently, the effects of noise deletion/addition on target deviation can be directly assessed. This upper bound provides actionable guidance for the joint optimization of training coverage and prediction strategies, and can undoubtedly be applied more broadly to the analysis and optimization of model performance.
@@ 6 Experimental Results
Using the feedforward neural network model listed in Appendix C, we conducted regression experiments on two bivariate root functions: $y_{1}=\sqrt{4x_{1}^{2}+3x_{2}^{2}}$ and $y_{2}=\sqrt{2x_{1}^{2}+x_{2}^{2}}$. Both the training and test datasets were specific. We evaluated the models’ prediction errors under various conditions, assessed the impact and computational cost of ethical and safety mechanisms, and explored the challenge of providing multiple types of explanations for the same output. These experiments serve to illustrate and validate the definitions and theorems presented in Section 5.
6.1 Experiments on the TVD upper bound under noiseless input and conservative prediction strategy
To evaluate the TVD upper bound estimation of Theorem 3 under noiseless input and conservative prediction strategy, as well as its relationship to model performance, we construct the following experimental setup. On the ${(x}_{1},x_{2})$ plane, we select a square region centered at (1.0,1.0) with side length 1, and uniformly sample 10,000 points as the training dataset. For the test dataset, we uniformly sample 10,000 points within a circular region centered at (1.5,1.0) with radius 0.5. The model learning rate is set to $\eta=0.01$. By uniformly sampling 10,000 points within the test region and computing the corresponding discrete points in the ${(y}_{1},y_{2})$ plane using the target function, we approximate the distribution $p_{q}(x)$ of the test dataset and its entropy function. Fig. 3 illustrates the experimental conditions and outcomes:
@@ 6.1 Experiments on the TVD upper bound under noiseless input and conservative prediction strategy
(c) For all test samples from the non-overlapping region, a conservative uniform prediction strategy was used. For almost all samples (except a few outliers), their RMSE is less than $TVD\approx\sqrt{((1-0.4951)\ln{((10000-4951)}/(1-0.4951))-2.9741)/2}\approx 0.9155$. This indicates that even in regression tasks, although TVD is not strictly equivalent to generalization error, it can still effectively capture the distributional characteristics of generalization error.
@@
Figure 3: Experimental Evaluation and Results of the TVD Upper Bound under Noiseless Input and Conservative Prediction Strategy. (a) The blue rectangular region represents the training dataset comprising 10,000 samples, while the red circular region represents the test dataset, also with 10,000 samples. Among these, 4,951 samples overlap with the training set. (b) For each of the training steps (10, 50, 100, 200, 500, 1000, 4000), the variation in root mean square error (RMSE) is shown for 10 randomly selected test samples from the training dataset. (c) In the non-overlapping region between the test and training datasets, the RMSEs for all 5,049 test samples are compared between the conservative uniform prediction strategy and the TVD upper bound. (d) For 10 randomly selected test samples from the non-overlapping region, the changes in RMSE are shown for training steps (10, 30, 300, 600, 1000, 4000, 7000), along with comparisons to both the TVD upper bound and the RMSE from the conservative uniform prediction.
6.2 Experimental Evaluation of the TVD Upper Bound under Noisy Input and Arbitrary Prediction Strategy
To evaluate Theorem 4’s estimation of the TVD upper bound under conditions of noisy input and the adoption of arbitrary prediction strategies in regions not covered by the training data, as well as itsrelationship to model performance, we utilize the same model instance, training dataset, and test dataset as in the previous experiment. We introduce noise into the input by locally adding or deleting noise information in the test dataset according to Definition 3, so that the model input contains noise. Meanwhile, in the regions not covered by the training data, we apply an arbitrary prediction strategy based on the joint distribution $p_{0}(x)$ formed by two independent normal distributions, $N\sim(0.6306,{0.4585}^{2})$ and $N\sim\left(0.4031,{0.3666}^{2}\right)$, to process and predict the input data. Fig. 3: (a) The training dataset $S_{ou}$ (selected within the blue square) and the test dataset $S_{oq}$ (selected within the red circle) are the same as those used in the experiment described in Section 6.1. The upper arc-shaped segment of $S_{oq}$ is shifted upward, and the vacated region, denoted as $S_{oq}^{-}$, is excluded from the test input. The shifted portion, $S_{oq}^{+}$, does not overlap with the original test dataset but partially overlaps with $S_{ou}$ ; it is added as new noisy test data. This structure satisfies the definition of noisy input, $\SoqSoq+\widetilde{S_{oq}}=S_{oq}\backslash$ $\cup S_{oq}^{+}$.
Figure 4: Experimental Evaluation and Results of the TVD Upper Bound under Noisy Input and Arbitrary Prediction Strategy. (a) The upper arc-shaped region of the test dataset is shifted upward, and the resulting vacant region is excluded from the test input. The shifted segment, which does not overlap with the original test dataset, is incorporated as noise into the new test dataset. (b) One hundred groups of samples are randomly selected from the noisy test dataset and processed by the model. The RMSE of the predictions are represented by 100 points along the blue line, while the red line indicates the TVD upper bound between the noisy input distribution and the model distribution. (c) In the case of $P_{l}\geq P_{a}$, 100 samples are randomly selected from the noiseless test dataset, transformed into the corresponding noisy data, and processed by the model. The RMSE of each prediction is represented by 100 points along the blue line, with the red line denoting the TVD upper bound between the noiseless input distribution and the model distribution. (d) In the case of $P_{l}<P_{a}$, 100 samples are randomly selected from the noiseless test dataset, transformed into the corresponding noisy data, and processed by the model. The RMSE of each prediction is represented by 100 points along the blue line, with the red line indicating the TVD upper bound between the noiseless input distribution and the model distribution.
(b) For the noisy test dataset comprising 11,034 samples, we compute the probability distribution for each test sample using the same method as in Section 6.1. The loss mass is $P_{l}=0.0744$, the addition mass is $P_{a}=0.0517$, the uncovered mass is $P_{uncov}=0.5091$, and the KL divergence is $D_{KL}\left(r\parallel p_{0}\right)=45.58$. This yields a TVD upper bound between the noisy input distribution and the model distribution of 3.312. By randomly selecting 100 groups of samples from the test dataset and inputting them into the model after 20 training steps, we find that the root mean square error (RMSE) of all outputs (marked by blue dots) is less than the TVD upper bound indicated by the red line. This validates the conclusion of Theorem 4, Equation (5.17).
(c) In the above example, $P_{l}=0.0744>0.0517=\gamma$. Under the same conditions, we randomly select 100 samples from the noiseless test dataset, convert them into corresponding noisy data as input to the model, and find that the RMSE of all outputs remains below the TVD upper bound of 3.4555. This validates the case where $P_{l}\geq P_{a}$ as given in Theorem 4, Equation (5.18).
(d) To examine the case where $P_{l}<P_{a}$ in Theorem 4, Equation (5.18), we construct a new noisy test dataset comprising 10,000 samples. In this case, $P_{l}=0.0744$, $P_{a}=0.1551$, $P_{uncov}=0.5530$, and $D_{KL}\left(r\parallel p_{0}\right)=40.88$, resulting in a TVD upper bound of 3.4325. Again, by randomly selecting 100 samples from the noiseless test dataset and converting them into noisy inputs for the model, the RMSE of all outputs remains below the TVD upper bound indicated by the red line, thereby confirming the corresponding conclusion provided by Theorem 4.
Through these experiments, we have comprehensively validated Theorem 4 regarding the TVD upper bound under noisy input and arbitrary prediction strategies. Furthermore, as illustrated in Fig. 4, there remains a noticeable gap between the RMSE of the model outputs and the TVD upper bound across various scenarios, indicating that the upper bound provided by Theorem 4 is still somewhat conservative. Further research may yield tighter error estimates under typical conditions.